A 15.01g sample of MgSO4*xH2O leave a residue of 6.01g upon heating. What is the value of x?
1
Expert's answer
2012-10-02T07:16:58-0400
15.01 6.01
MgSO4*xH2O ------> MgSO4 + xH2O
120+18X & 120

15.01*120/6.01=120+18x
300=120 + 18X
180=18x
& x=10